3 A random of sample of size 25 is taken with replacement from a population with 121.4 and 50.5. 4. A random sample of 20 independent observations is taken from a population with p = 23.8 and a = 5.

1
Expert's answer
2022-06-07T00:16:09-0400

3.


"\mu_{\bar{X}}=\mu=121.4"

"\sigma^2_{\bar{X}}=\sigma^2/n=50.5/25=2.02"

"\sigma_{\bar{X}}=\sqrt{\sigma^2_{\bar{X}}}=\sqrt{2.02}\approx1.421267"

4.


"\mu_{\bar{X}}=\mu=23.8"

"\sigma^2_{\bar{X}}=\sigma^2/n=5^2/20=1.25"

"\sigma_{\bar{X}}=\sqrt{\sigma^2_{\bar{X}}}=\sqrt{1.25}\approx1.118034"